The rate constant for a first order reaction is 0.20 M-1s-1. The time it takes for...

  1. The rate constant for a first order reaction is 0.20 M-1s-1. The time it takes for 0.50 M of reactant to halve is________ s

    1.2

    2.5

    3.5

    10
Answer #1

pale constone = K = 0.20 m!5 it is a second order kinetic whose half life is given half life = t1/2 = K(A). TA). → Initial c